# Mixed-Strategy Nash Equilibrium for Crowd Navigation

### v2

@@ VI-C Sampling-based crowd navigation
VI-D Algorithm complexity
We analyze the computational time complexity of a single iteration within the sampling-based algorithm. Given $T$ time steps as planning horizon, $N$ agents, and $M$ samples for each agent, the complexity of computing (72) for rejection-based sampling and (73) for importance sampling is $O(TMN)$. For each iteration of the Bayesian updating scheme, (72) and (73) need to be computed for all the agents, which leads to an overall complexity of $O(KMN^{2})$ for one iteration. On the other hand, the convergence rate of the iterative Bayesian updating scheme depends on the nominal mixed-strategies, which are unknown a priori. However, in Figure 6 we can see that the number of iterations required for the algorithm to converge does not depend on the number of players. Thus, our sampling-based crowd navigation in practice has a quadratic computation complexity of $O(KMN^{2})$ with respect to the number of agents. Note that, given $N$ agents, $O(N^{2})$ is the minimal complexity required for reasoning over all possible two-agent interaction pairs. The computation complexity of our algorithm is lower than the state-of-the-art dynamic game solver ALGAMES [60], which has a computation complexity of $O(MN^{3})$. Furthermore, the computation of sampling weights in (72) and (73) are highly parallelizable, which further reduces the computation time in practice.
VII Evaluation
VII-A Overview of evaluation
@@
We evaluate our BRNE navigation model, as well as baseline methods, in two categories of tasks: (1) multi-agent navigation (Section VII-B); (2) crowd navigation (Section VII-C, Section VII-D). In the multi-agent navigation experiments, the tested method is asked to plan paths for all the agents, given a set of start locations and goal locations. Multi-agent navigation experiments represent the most ideal situation of crowd navigation, where all agents’ behaviors follow the assumption of the tested model. We evaluate multi-agent navigation performance as a lower bound of the algorithm’s ability to infer efficient and safe joint paths. However, for real-world navigation in human crowds, pedestrians do not necessarily follow the behavior assumed by the planning algorithm. Therefore, we conduct extensive crowd navigation experiments to test different algorithms’ performance in realistic crowd navigation scenarios. Furthermore, we conduct two kinds of crowd navigation experiments in different situations. The first crowd navigation evaluation (Section VII-C) places all the agents in a tight space where collision avoidance is necessary and simulates pedestrian behavior using a decentralized collision avoidance algorithm. The purpose of this evaluation is to stress test the robot’s ability to infer and adapt to cooperative but unknown joint navigation strategies from other agents. The second crowd navigation (Section VII-D) aims to benchmark the algorithm performance in realistic crowd navigation tasks. There, we conduct the evaluation in a state-of-the-art social navigation benchmark, which consists of real-world environmental geometry and real human navigation trajectories. The robot has to navigate in unstructured spaces while interacting with human pedestrians in non-controlled environments. For each experiment, we evaluate the algorithm performance from two categories of metrics: safety and navigation efficiency. Details of the specific metrics are provided in the corresponding subsections. We will release the code of our implementation.
VII-B Multi-agent navigation evaluation
@@
For multi-agent navigation, we randomly place different numbers of agents on a circle with the radius being 3 meters. The goal of each agent is to reach the other side of the circle. This design means the most efficient path for each agent is to cross the center of the circle, which makes it necessary for all agents to compromise efficiency in exchange of safety. When randomizing the initial agent locations, we uniformly sample locations on the circle, but also make sure no two agents’ initial locations are within 0.6 meters of each other—we assume each agent is circular-shaped with a radius of 0.3 meters. We vary the number of agents from 4 to 8 and conduct 100 navigation trials for each number of agents. We specify the desired velocity for all agents and all baselines as $1.2m/s$. We test our method BRNE with both sampling schemes, we denote the rejection-sampling-based implementation as BRNE(RS) and the importance-sampling-based implementation as BRNE(IS).
@@
For safety, we measure safety distance and collision rate. The safety distance is defined as the closest distance between any two agents during the whole navigation task. We consider a navigation trial involves collision if the safety distance of the trial is below 0.6 meters, twice the body radius of each agent. For efficiency, we measure the maximum path length among all agents in each trial.
Results
The evaluation results are presented in Table I(b) (safety) and Table II (navigation efficiency). Note that ORCA guarantees collision-free paths, thus it has $0\%$ collision rate across all experiments. On the other hand, our method BRNE generates joint navigation strategies with larger safety distance, better efficiency, and lower disparity. Even though BRNE strategies are not collision free, the collision rate is still minimal. Furthermore, as shown in Figure 5, ORCA is symmetry-sensitive, which leads to highly unnatural joint strategies when agents are unevenly distributed. Lastly, the learning-based baseline CADRL performs poorly across all metrics and exhibits highly non-smooth trajectories in Figure 5. Note that, the CADRL policy is trained using the toolbox provided in [52] with pedestrians simulated with ORCA. Thus, the trained CADRL policy is a collision avoidance policy in crowds. We test the CADRL policy in this multi-agent navigation benchmark to verify if the policy suffices as a coupled-prediction-and-planning method, instead of a dynamic object avoidance method. A detailed demonstration of the multi-agent navigation experiment results can be found in Extension 1.
Figure 6: Convergence of mixed-strategy across different numbers of agents in the multi-agent navigation experiments.
@@ Results
Computation time
We use the multi-agent navigation benchmark to also evaluate how the number of agents and the number of samples influence the computation time of BRNE. We fixed the planning horizon of BRNE at 50 time steps (5 seconds), then we vary the number of agents from 4 to 8 and the number of samples per agent from 100 to 500. The tested BRNE implementation uses Numba for parallelization and runs on a laptop with Intel Core i9-12900K Processor (20 threads). The results are shown in Figure 7. We can see that when the algorithm is accelerated through parallelization in practice, the computation complexity is close to linear with respect to the number of agents. In Table III, we further compare the computation speed of BRNE with two state-of-the-art dynamic game solvers, ALGAMES [60] and iLQGames [59]. We use 100 samples per agent for BRNE and use the results reported in the “intersection” experiment from [60]. Throughout this test, we use the importance-sampling-based BRNE implementation.
TABLE III: Comparison of computation time.
@@
Convergence test
We use the multi-agent navigation benchmark to further test the convergence of the BRNE algorithm. In Figure 6, we show the convergence of BRNE with the number of agents varying from 4 to 8. We can see the algorithm consistently converges within 10 iterations and the convergence rates are similar across different numbers of agents. For better numerical accuracy, we use the rejection-sampling-based implementation for the convergence test.
VII-C Simulated crowd navigation evaluation
@@
The geometrical layout of the simulated crowd navigation experiment is similar to the multi-agent navigation experiment, where we place agents across a circle with a radius of 3 meters and each agent’s task is to reach the other side of the circle. The difference is, in the crowd navigation experiment we only have control over one agent (the robot), while other agents are cooperative but make independent decisions. We use ORCA to simulate pedestrians since it is the most commonly used decentralized collision avoidance method. We test each navigation method with $5$ simulated pedestrians. The desired velocity for each agent is set as $1.2m/s$. The body radius of each agent is $0.3m$, thus the distance threshold for a collision is $0.6m$. We randomly initialize agent locations for 100 trials, with the same randomization scheme in the multi-agent navigation experiments. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed-strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
@@
For safety, we measure the same safety distance and collision rate metrics. For navigation efficiency, we measure the time-to-goal and path length.
@@ Results
In Table IV and Figure 8, we show the safety and navigation efficiency results of the simulated crowd navigation experiments. In Figure 9, we show frames of our method from the simulated tests. We can see that our method BRNE is competitive on both navigation safety and efficiency with the state-of-the-art reinforcement learning method NCE. Both NCE and CADRL are tested in the same simulation environment where it is trained, while our method BRNE requires no training. CADRL performs poorly in this benchmark, exhibiting high collision rate and low navigation efficiency. Furthermore, CADRL fails to reach the navigation goal and we have to terminate the trial after a certain amount of time, causing the zero standard deviation in the time-to-goal metric. The performance of CADRL could be due to the difference between the testing environment, which is provided by [54], and the training environment of CADRL. A detailed demonstration of the experimental results of our method BRNE in the simulated crowd navigation benchmark can be found in Extension 1.
VII-D crowd navigation in human datasets
@@
We evaluate crowd navigation with prerecorded human pedestrian behaviors within the same geometrical spaces of recordings. The evaluation is conducted in SocNavBench [78], a state-of-the-art, high-fidelity social navigation benchmarking framework. SocNavBench contains human pedestrian data from two of the most commonly used pedestrian datasets ETH [79] and UCY [71]. A set of 33 curated episodes are extracted from the original datasets for benchmarking, which assemble a set of highly interactive test trials. The curated episodes contain an average of 44 pedestrians per episode. The distance between the goal location and the start location for the robot makes sure the navigation task can be finished within 25 seconds when the robot does not perform collision avoidance and navigates with the maximum permitted velocity of $1.2m/s$. An occupancy grid map of the real-world space from the original dataset is provided for each episode, the robot has access to the map a priori and can only navigate in the free space. Note that, even though the pedestrian agents in SocNavBench are non-reactive given the pre-recorded human behavior, the robot agent still assumes the pedestrian agents are real-world reactive agents. We refer readers to [78] for more details about the implementation of the benchmark, as well as the discussion on the choice of using pre-recorded human behavior. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed-strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
@@
We compare our method BRNE with the four baselines reported in the original paper of SocNavBench [78], which are (1) the social force model (SF[7]; (2) ORCA; (3) CADRL; and (4) a pedestrian-unaware baseline with static obstacle avoidance capability (Baseline). Note that ORCA and CADRL do not have static obstacle avoidance capability, thus they use Baseline as a meta-planner. Similarly, we implement our meta-planner with static obstacle avoidance capability for BRNE based on a constant velocity model.
@@
We use the four main metrics based on the original metrics provided by the benchmark: (1) The total number of pedestrian collisions; (2) The total number of freezing behaviors, where the robot fails to reach the goal within 60 seconds or runs into environmental obstacles; (3) Path length; (4) Time to goal. Full results with complete benchmark metrics are provided in the appendix.
@@
Table V shows results of the human dataset crowd navigation experiments. Figure 10 shows representative snapshots of the benchmark tests. Both our method (BRNE) and the social force model (SF) significantly outperform the rest of the methods on navigation safety. Our method shares the minimal number of collisions with the social force model while being the only method with zero freezing behavior and exhibiting better path efficiency compared with the social force model. More importantly, compared to our meta-planner baseline Meta-planner(Ours), whose path is used by BRNE as the nominal strategy, BRNE significantly improves navigation safety, reducing the number of pedestrian collisions from 37 to 1, while sharing nearly identical navigation efficiency as the baseline. Further, as discussed in the next subsection, this one collision incident of our method is due to the pre-processing of the dataset instead of our method’s performance. These experimental results show that, when interacting with real human crowds in unstructured environments, our method is able to improve an existing navigation method’s safety to be near-perfect with minimal impact on navigation efficiency. This indicates that our method reaches human-level crowd navigation performance, on both safety and navigation efficiency, on top of a meta-planner, and also shows great potential for deploying our method on other existing navigation pipelines. A detailed demonstration of the human dataset benchmark can be found in Extension 1.
@@
There is one trial where the robot collides with a pedestrian, here we provide an analysis of the cause, with the snapshots of the collision shown in Figure 11. As we can see, the sudden appearance of a pedestrian right in front of the robot, which is an artifact of the pre-processing of the dataset, leaves the robot with no space and no time to react. Therefore, this collision incident does not reflect our method’s navigation safety performance and can be avoided in practice. A detailed demonstration of the failure trial can be found in Extension 1.
@@
While we leave testing our method in real-world crowds for future work, due to the time and financial cost of conducting such experiments, we want to discuss the potential gap of sim-to-real transfer based on the human dataset benchmark results. First, since the robot is observing real-world human behavior in the benchmark, the benchmark results can at least verify which method will fail in the real world—a method that is dangerous or inefficient in the benchmark will likely remain dangerous or inefficient in the real world. Second, one main challenge of sim-to-real transfer in crowd navigation is the misalignment of human behavior assumptions. In our case, by inferring mixed-strategy Nash equilibrium for navigation, the robot assumes the humans are cooperative, which is not necessarily true in reality. However, the human dataset benchmark can serve as a preliminary verification of the influence of the misaligned human behavior assumption. Recall that, even though human agents are not reactive in the benchmark, the robot still believes it is interacting with real humans and is indeed observing real human behavior. The numerical efficiency of our method enables fast replanning during navigation, allowing the robot to constantly adjust its plan based on the latest observation, including observations that humans do not follow the assumed cooperative behavior. Furthermore, the robot makes navigation decisions based on mixed-strategy Nash equilibrium, it does not anticipate exact human behavior but instead a distribution of likely human behaviors, which already takes into account unexpected human reactions. Thus, the fast replanning capability and mixed-strategy-based modeling make the performance of our method in the human dataset benchmark more transferable in the real world. Lastly, our method is essentially built on top of an existing meta-planner. The modular design of our algorithm enhances the deployment flexibility, which lowers the gap of sim-to-real transfer. For example, our method can be integrated into existing robot navigation frameworks such as the navigation stack from the Robot Operating System. This is particularly important for crowd navigation in unstructured real-world environments, which most existing crowd navigation methods and all game-theoretic planning methods do not address. Lastly, even though out of the scope of this paper, we want to point out one important aspect of the sim-to-transfer that is not addressed in the benchmark: the perception aspect of crowd navigation. The benchmark assumes near-perfect robot perception of pedestrian states, but robust perception for both robot localization and pedestrian tracking remains an open challenge and could influence the real-world deployment of our method.
VIII Conclusion

### v4

@@ V Evaluation
Figure 6: Convergence of mixed strategy in our algorithm across different numbers of agents in the multi-agent navigation experiments.
@@
V-A Overview of evaluation
@@
We evaluate our BRNE navigation method, as well as other non-learning and learning-based crowd navigation methods, in two categories of tasks: (1) multi-agent navigation (Section V-B); (2) crowd navigation (Section V-C, Section V-D). In the multi-agent navigation experiment, the tested methods are asked to plan paths for all the agents, given a set of start and goal locations. The multi-agent navigation experiment represents the most ideal situation of crowd navigation, where all agents’ behaviors follow the behavioral model assumed by the planning algorithm. The results of the multi-agent navigation experiment do not necessarily reflect an algorithm’s crowd navigation performance with humans, but serve as an upper bound of the algorithms’ ability to infer safe and efficient joint paths—if a method cannot infer safe and efficient joint paths even when all the agents follow the assumed behavioral model, then the method is unlikely to generate safe and efficient navigation actions with real humans who do not necessarily follow the assumed behavioral model. To evaluate the crowd navigation performance of the tested methods, especially when pedestrians do not follow the behavior assumed by the planning algorithm, we conduct extensive crowd navigation benchmark experiments with two different kinds of experiment designs. The first crowd navigation benchmark experiment (Section V-C) places all the agents in a tight space where collision avoidance is necessary for all the agents. The pedestrian behavior is simulated using a decentralized collision avoidance algorithm ORCA [8]. The purpose of this evaluation is to stress test the robot’s ability to infer and adapt to cooperative but unknown joint navigation strategies from other agents. The second crowd navigation benchmark experiment (Section V-D) aims to benchmark the algorithms’ performance in realistic crowd navigation scenarios. We evaluate the performance using a state-of-the-art crowd navigation benchmark [81], which consists of real-world environmental geometry and real human navigation trajectories collected from several datasets. The robot has to navigate in unstructured spaces while interacting with human pedestrians. For each experiment, we evaluate the algorithm performance from two categories of metrics: safety and navigation efficiency. Details of the specific metrics are provided in the corresponding subsections. Lastly, we implement the algorithm on an untethered quadruped robot with fully onboard perception and computation to demonstrate our algorithm’s practicality with real humans. Videos of the benchmark experiments and hardware demonstration are included in the project website: https://sites.google.com/view/brne-crowdnav. We will also release the code and parameters of our implementation.
V-B Multi-agent navigation evaluation
@@
For multi-agent navigation, we randomly place different numbers of agents on a circle with a radius of 3 meters. The goal of each agent is to reach the other side of the circle. This design means the most efficient path for each agent is to follow a straight line across the center of the circle, which makes it necessary for all agents to compromise efficiency in exchange for safety. When randomizing the initial agent locations, we uniformly sample locations on the circle, but also make sure no two agents’ initial locations are within 0.6 meters of each other—we assume each agent is circular-shaped with a radius of 0.3 meters. We vary the number of agents from 4 to 8 and conduct 100 navigation trials for each number of agents. We specify the desired velocity for all agents and all baselines as $1.2m/s$. We test our method BRNE with both an importance-sampling implementation and a rejection-sampling implementation, we denote the rejection-sampling implementation as BRNE(RS) and the importance-sampling implementation as BRNE(IS).
@@
For safety, we measure safety distance and collision rate. The safety distance is defined as the closest distance between any two agents during the whole navigation task. We consider a navigation trial to involve a collision if the safety distance of the trial is below 0.6 meters, twice the body radius of each agent. For the navigation efficiency, we measure the maximum path length among all agents in each trial.
@@
TABLE III: Computation time of game-theoretic methods.
@@
Results
The evaluation results are presented in Table I(b) (safety) and Table II (navigation efficiency). Note that ORCA guarantees collision-free paths, thus it has $0\%$ collision rate across all experiments. On the other hand, our method BRNE generates joint navigation strategies with larger safety distance, better efficiency, and lower disparity. Even though BRNE strategies are not collision-free, the collision rate is still minimal. Since the multi-agent navigation benchmark assumes all agents follow the same behavioral model, having a $0\%$ collision rate in this benchmark does not necessarily mean collision-free with humans. More importantly, as shown in Figure 5, the collision-free guarantee of ORCA could significantly compromise the path quality, leading to highly unnatural joint strategies when agents are unevenly distributed. This indicates that, when used as a coupled prediction and planning framework on the robot, ORCA could plan the robot’s path based on highly unrealistic expectations of human reactions. Lastly, the learning-based baseline CADRL performs poorly across all metrics and exhibits highly non-smooth trajectories as shown in Figure 5. Note that the CADRL policy is trained using the toolbox provided in [52] with pedestrians simulated with ORCA. Thus, the trained CADRL policy is a collision avoidance policy in crowds. We test the CADRL policy in this multi-agent navigation benchmark to verify if the policy is a coupled prediction and planning method, instead of a dynamic object avoidance method. Videos of the multi-agent navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
@@ Results
Computation time of BRNE
We use the multi-agent navigation benchmark to evaluate how the number of agents and samples affect the computation time of BRNE. We fixed the planning horizon of BRNE at 50 time steps (5 seconds), then we vary the number of agents from 4 to 8 and the number of samples per agent from 100 to 500. The tested BRNE implementation uses Numba for parallelization and runs on a laptop with Intel Core i9-12900K Processor (20 threads). The results are shown in Figure 7. We can see that when the algorithm is accelerated through parallelization in practice, the computation complexity is close to linear with respect to the number of agents. Since the scalability with the number of agents is often a computation bottleneck for game-theoretic planners, in Table III, we further compare the computation speed of BRNE with two state-of-the-art dynamic game solvers, ALGAMES [60] and iLQGames [59]. We use 100 samples per agent for BRNE and use the results reported in the “intersection” experiment from [60], as it is the most similar to the crowd navigation setting in this test. We use the importance-sampling-based BRNE implementation throughout this test. In Table III, we can see that BRNE is at least one order of magnitude faster than ALGAMES and iLQGames and is the only method with sufficient computation time for real-time planning with 4 agents. ALGAMES and iLQGames are general-purpose game-theoretic planners that can be applied to applications other than crowd navigation, while BRNE is designed specifically for crowd navigation. BRNE explicitly benefits from its narrow scope for better computation efficiency in dense crowds.
Convergence test
We use the multi-agent navigation benchmark to further test the convergence of the BRNE algorithm. In Figure 6, we show the convergence of BRNE with the number of agents varying from 4 to 8. We can see the algorithm consistently converges within 10 iterations and the convergence rates are similar across different numbers of agents. We use the rejection-sampling-based implementation for the convergence test for better numerical accuracy.
V-C Simulated crowd navigation evaluation
@@
The geometrical layout of the simulated crowd navigation experiment is similar to the multi-agent navigation experiment, where we place agents across a circle with a radius of 3 meters and each agent’s task is to reach the other side of the circle. The difference in this benchmark is that we only have control over one agent (the robot), while other agents are cooperative but make independent decisions. We use ORCA to simulate pedestrians since it is the most commonly used decentralized collision avoidance method. We test each navigation method with $5$ simulated pedestrians. The desired velocity for each agent is set as $1.2m/s$. The body radius of each agent is $0.3m$ and the distance threshold for a collision is $0.6m$. We randomly initialize agent locations for 100 trials, with the same randomization scheme in the multi-agent navigation experiments. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
@@
For safety, we measure the same safety distance and collision rate metrics. For navigation efficiency, we measure the time-to-goal and path length of the robot.
@@ Results
In Table IV and Figure 8, we show the safety and navigation efficiency results. In Figure 9, we show representative frames of our method from the simulated tests. We can see that our method BRNE is competitive on both navigation safety and efficiency with the state-of-the-art reinforcement learning method NCE. Both NCE and CADRL are tested in the same simulation environment where it is trained, while our method BRNE requires no training. CADRL performs poorly in this benchmark, exhibiting a high collision rate and low navigation efficiency. Furthermore, CADRL fails to reach the navigation goal and we have to terminate the trial after a certain amount of time, causing the low standard deviation in the time-to-goal metric. The performance of CADRL could be due to the difference between the testing environment provided by [54] and the training environment of CADRL. Videos of the simulated crowd navigation experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
V-D Crowd navigation in human datasets
@@
We evaluate crowd navigation with prerecorded human pedestrian behaviors within the same geometrical spaces of recordings. The evaluation is conducted in SocNavBench [81], a state-of-the-art, high-fidelity crowd navigation benchmark framework. SocNavBench contains human pedestrian data from two of the most commonly used pedestrian datasets ETH [82] and UCY [83]. A set of 33 curated episodes are extracted from the original datasets for benchmarking, which assemble a set of highly interactive test trials. The curated episodes contain an average of 44 pedestrians per episode. The distance between the goal location and the start location for the robot ensures that the navigation task can be finished within 25 seconds if the robot does not avoid collision and navigates with the maximum permitted velocity of $1.2m/s$. An occupancy grid map of the real-world space is provided for each episode, the robot has access to the map a priori and can only navigate in the free space. Note that, even though the pedestrian agents in SocNavBench are non-reactive given the pre-recorded human behavior, the robot agent still assumes the pedestrian agents are real-world reactive humans. We refer readers to [81] for more details about the benchmark, as well as the discussion on the choice of using pre-recorded human behavior. In this experiment, BRNE is implemented with importance sampling as a model predictive planner, updating its inference of mixed strategy Nash equilibrium based on the latest observation of other agents’ positions and velocities.
@@
We compare our method BRNE with the four baselines and the corresponding results reported in the original paper of SocNavBench [81]. They are (1) the social force model (SF[7]; (2) ORCA; (3) CADRL; and (4) a pedestrian-unaware baseline provided by the benchmark with static obstacle avoidance capability (Baseline). Note that ORCA and CADRL do not have static obstacle avoidance capability. Thus, they use Baseline as a meta-planner to avoid static obstacles in the benchmark. We tried to use Baseline directly as the meta-planner to generate the nominal mixed strategy for BRNE. However, even though Baseline is sufficient for BRNE to avoid static obstacles, it is not sufficient for BRNE to consistently reach the goal—recall that BRNE requires a strong goal-oriented nominal mixed strategy. Therefore, we slightly modified Baseline to keep its static obstacle avoidance module but improve its goal-reaching capability. We name this modified baseline planner Meta-Planner and report its performance as well. We will release the implementation of both BRNE and Meta-Planner in this benchmark.
Figure 12: Snapshots from the real-world demonstration with 4 pedestrians. The specification of this demonstration mimics the simulation tests shown in Figure 9. Videos of the demonstrations are included in the project website: https://sites.google.com/view/brne-crowdnav.
Figure 13: Snapshots from the real-world demonstration with 10 pedestrians. The specification of this demonstration mimics the benchmark tests shown in Figure 10. Videos of the demonstrations are included in the project website: https://sites.google.com/view/brne-crowdnav.
@@
We use the four main metrics: (1) The total number of pedestrian collisions; (2) The total number of freezing behaviors, where the robot fails to reach the goal within 60 seconds or runs into environmental obstacles; (3) Path length; (4) Time to goal. Full results with complete benchmark metrics are provided in the appendix.
@@ Results
Table V shows results of the human dataset crowd navigation experiments. Figure 10 shows representative snapshots of the benchmark tests. Both our method (BRNE) and the social force model (SF) significantly outperform the rest of the methods on navigation safety. Our method shares the minimal number of collisions with the social force model while being the only method with zero freezing behavior and exhibiting better path efficiency compared with the social force model. More importantly, compared to our meta-planner baseline Meta-planner(Ours), whose path is used by BRNE to generate the nominal mixed strategy, BRNE significantly improves navigation safety, reducing the number of pedestrian collisions from 37 to 1, while sharing nearly identical navigation efficiency as the baseline. As discussed in the next subsection, this one collision incident of our method is due to the pre-processing of the dataset instead of our method’s performance. These experimental results show that, when interacting with real human crowds in unstructured environments, our method can improve an existing navigation method’s safety to be near-perfect with minimal impact on navigation efficiency. This indicates that our method reaches human-level crowd navigation performance with a meta-planner on safety and navigation efficiency. The results also show great potential for deploying our method on other navigation pipelines. Videos of the human dataset experiments are included in our project website: https://sites.google.com/view/brne-crowdnav.
@@
There is one trial where the robot collides with a pedestrian, here we provide an analysis of the cause, with the snapshots of the collision shown in Figure 11. As we can see, the sudden appearance of a pedestrian right in front of the robot, which is an artifact of the pre-processing of the dataset, leaves the robot with no space and no time to react. Therefore, this collision incident does not reflect our method’s navigation safety performance and can be avoided in practice. The video of the failure trial is included in our project website: https://sites.google.com/view/brne-crowdnav.
@@
While we will demonstrate our BRNE algorithm on robot hardware with real-world humans in the next section, we leave benchmarking our method in real-world crowds for future work, due to the time and financial cost of conducting such benchmarks. Here, we want to discuss the potential gap of sim-to-real transfer based on the human dataset benchmark results. First, since the robot is observing real-world human behavior in the benchmark, the benchmark results can at least verify which method will fail in the real world—a method that is dangerous or inefficient in the benchmark will likely remain dangerous or inefficient in the real world. Second, one main challenge of sim-to-real transfer in crowd navigation is the misalignment of human behavior assumptions. In our case, by inferring mixed strategy Nash equilibrium for navigation, the robot assumes the humans are cooperative, which is not necessarily true. However, the human dataset benchmark can serve as a preliminary verification of the influence of the misaligned human behavior assumption. Recall that, even though human agents are not reactive in the benchmark, the robot still believes it is interacting with real humans and observing real human behavior. The numerical efficiency of our method enables fast replanning during navigation, allowing the robot to constantly adjust its plan based on the latest observation, including observations that humans do not follow the assumed cooperative behavior. Furthermore, the robot makes navigation decisions based on mixed strategies—it does not anticipate exact human behavior but instead a distribution of likely human behaviors, which already considers unexpected human reactions. Thus, the fast replanning capability and mixed strategy-based modeling make the performance of our method in the human dataset benchmark more transferable in the real world. Lastly, our method is essentially built on top of an existing meta-planner provided by the benchmark. The modular design of our algorithm enhances the deployment flexibility, which lowers the gap of sim-to-real transfer. For example, our method can be integrated into existing robot navigation frameworks such as the navigation stack from the Robot Operating System [84, 85]. This is particularly important for crowd navigation in unstructured real-world environments, which most existing crowd navigation methods and all game-theoretic planning methods do not address. Lastly, even though out of the scope of this paper, we want to point out one important aspect of the sim-to-transfer that is not addressed in the benchmark: the perception aspect of crowd navigation. The benchmark assumes near-perfect robot perception of pedestrian states, but robust perception for both robot localization and pedestrian tracking remains an open challenge and could influence the real-world deployment of our method.
V-E Real-World Hardware Demonstration
We demonstrate the practicality of our algorithm in the real world on an untethered robot quadruped with fully on-board perception and computation. The robot successfully completed 10 real-time crowd navigation tasks without collision or freezing behavior, with the number of real human pedestrians varying from 4 to 10. Figure 12 and Figure 13 show snapshots from two representative demonstrations. We include videos of all 10 demonstrations on our project website: https://sites.google.com/view/brne-crowdnav.
Note that the demonstrations are only for verifying the algorithm’s practicality under the conditions and limitations of the real-world environment and robot hardware. They are not intended to be studies for benchmarking the algorithm’s performance in real-world human crowds, as such studies are out of the scope of this paper and would require implementing other algorithms in hardware for comparison.
System Specification
We use a Unitree Go1 EDU quadruped robot as the demonstration platform. Even though the robot has holonomic dynamics, we constrain its dynamics to be a differential-drive vehicle to demonstrate the compatibility of our algorithm with a broader range of mobile robots. We use a ZED 2i camera mounted onboard as the perception module. The perception module localizes the robot using visual inertia odometry at 50 Hz and tracks pedestrian position and velocity at 15 Hz. All the computation, including perception and planning with BRNE, is processed entirely by an Nvidia Jetson AGX Orin embedded computer mounted on the robot.
Software Implementation
We implement our BRNE algorithm within the Isaac Robot Operation System (ROS) framework to communicate with the perception module and the robot’s low-level controller. We implement the algorithm in Python and used the PyTorch library to accelerate the weight updating step of the algorithm using the low-power GPU of the onboard computer. To maintain the real-time computation speed, we limited the maximum number of agents processed by the algorithm to 5, which allows the algorithm to control the robot at 10 Hz and replan at 5 Hz, with a planning horizon of 2 seconds and 200 trajectory samples per agent. We will release the hardware design and software implementation alongside the algorithm parameters.
Demonstration Design
The demonstrations are conducted in a $3m$ wide and $9m$ long space inside an indoor atrium. The demonstrations are designed to mimic the specifications of the crowd navigation tests with ORCA agents (such as the ones shown in Figure 9) and with pedestrian datasets (such as the ones shown in Figure 10). The number of pedestrians varies from $4$ to $10$. Our demonstrations follow common design principles in other crowd navigation works, such as [43] and [45]. To mimic the test with ORCA agents, pedestrians were instructed to go to the designated goals in a circle. To mimic the pedestrian dataset benchmark, pedestrians were instructed to move freely in the space. Across all demonstrations, the pedestrians are instructed to “move with normal walking speed and treat the robot as a walking person or dog”. The navigation task for the robot is to move safely and efficiently from one side of the room to the other, with a traveling distance varying between $6m$ to $8m$. The robot moves at a nominal speed of $0.5m/s$.
VI Conclusion